OpenStudy (anonymous):

a surveyor, walking on a level ground at 10 km/hr in the eastward direction, sighted the foot of a tall narra tree in the direction E 12 degrees North. after 30mins, the same tree now bears East 48 degrees North . how far was he from the tree during the first sighting?

OpenStudy (anonymous):

That's a fast walk!

OpenStudy (anonymous):

could you help me to answer this question ?

OpenStudy (anonymous):

Yes, working on it now...

OpenStudy (anonymous):

Ok, I think the situation looks like this: 

OpenStudy (anonymous):

yah that is the exact illustration here in my book .

OpenStudy (anonymous):

You have angle-side-angle information about a triangle; you can easily find the third angle, then use a combination of law-of-sines and law-of-cosines to solve.

OpenStudy (anonymous):

the answer is 36 degrees?

OpenStudy (anonymous):

That's your third angle, yes. Do you know the law of sines and law of cosines?

OpenStudy (anonymous):

I think law of sines is all you need for this one. \[\frac{a}{\sin(A)}=\frac{b}{\sin(B)}\]

OpenStudy (anonymous):

I really can't remember how to use the law of sine and low of cosine.

OpenStudy (anonymous):

what about the law of cosine?

OpenStudy (anonymous):

Law of cosines is better when you have a side-angle-side situation. You can still use it for redundancy to check you answer.

OpenStudy (anonymous):



OpenStudy (anonymous):

so the answer is 7.01 km?

OpenStudy (anonymous):

Yep, that looks good to me.

OpenStudy (anonymous):

thank you I will post another question tomorrow I hope you find time to answer my questions.

OpenStudy (anonymous):

Just in case, I'm going to check it using law of cosines. Sometimes with obtuse triangles, sines can give ambiguous answers.

OpenStudy (anonymous):

@exeal Yeah, ran into some obtuse triangle trouble. Law of cosines gives a different answer. Both equations agree on d=6.32km. 7.01 is erroneous.

OpenStudy (anonymous):

so I will use the law of cosine. Thank you again

OpenStudy (anonymous):

I always use both to be on the safe side.

OpenStudy (anonymous):

You're welcome.